OpenStudy (anonymous):

Okay, so im starting to get the hang of this, but im not sure. 2(3x-4)=8x+2. Do i distribute whats in the ( ) then try and solve?

OpenStudy (aroub):

yes

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

that would be a fine first step :)

OpenStudy (anonymous):

first distribute and get your x's on a side non x's on the other

OpenStudy (anonymous):

So it would be 6x-6=8x+2? Or did i do that totally wrong..?

OpenStudy (anonymous):

check your multiplication 2*4=?

OpenStudy (amistre64):

distribution is just mulitplication: 3x-4 2 ----- 6x -8 ; so you was close

OpenStudy (anonymous):

So then i could do.. 6x-8=8x+2 6x=8x+10 -2x=10? Right,?

OpenStudy (amistre64):

so far so good, youre doing great :)

OpenStudy (anonymous):

yes now you need to get x by it's self

OpenStudy (anonymous):

So the answer would be x=-5? Or did i do my division wrong...?

OpenStudy (amistre64):

-5 is correct, good job!!

OpenStudy (anonymous):

nope you got it congrats